# Magnon Correlation Enables Spin Injection, Dephasing, and Transport in Canted Antiferromagnets

### v1

@@ Abstract
Thermal and electrical injection and transport of magnon spins in magnetic insulators is conventionally understood by the non-equilibrium population of magnons. However, this view is challenged by several recent experiments in noncollinear antiferromagnets, which urge a thorough theoretical investigation at the fundamental level. We find that the magnon spin in antiferromagnets is described by a matrix, so even when the diagonal terms—spins carried by population—vanish, the off-diagonal correlations transmit magnon spins. Our quantum theory shows that a net spin-flip of electrons in adjacent conductors creates quantum coherence between magnon states, which transports magnon spins in canted antiferromagnets, even without a definite phase difference between magnon modes in the incoherent process. It reveals that the pumped magnon correlation is not conserved due to an intrinsic spin torque, which causes dephasing and strong spatial spin oscillations during transport; both are enhanced by magnetic fields. Spin transfer to proximity conductors can cause extrinsic dephasing, which suppresses spin oscillations and thereby gates spin transport.
Introduction.—Antiferromagnets (AFMs) hold significant promise for spintronic applications due to their terahertz-frequency dynamics, immunity to external magnetic disorder, and ultrafast responses [1, 2, 3, 4, 5]. Spin angular momentum in collinear easy-axis AFMs is carried by circularly polarized magnon eigenmodes, which shows low dissipation in insulators with efficient long-range spin transport [6, 8, 9, 7, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 21, 22, 20, 23, 24]. In contrast, magnon modes in easy-plane AFMs are linearly polarized and are believed to carry no spins [3, 25, 26, 27] and the modes in canted AFMs have no spin projection along the Néel vector; both appear to be less likely to support spin currents [3, 16, 28, 29]. Experiments on canted easy-plane hematite ($\alpha$-Fe2O3) deny this intuition by observing efficient long-range magnon spin transport and a field-tunable Hanle effect [30, 31, 32]. The observation is interpreted as an interference of two linearly polarized modes, occasionally mimicking a circular polarization or pseudospin [5, 33, 34, 35].
@@
In this Letter, we demonstrate in the quantum description that the pseudospin in the classical description is a “real spin” carried by magnons by showing that the magnon spin is described by a matrix. When the spin carried by magnon population vanishes due to specific interactions, as shown in Fig. 1(a), the spin is solely carried by correlation ${\cal A}$, represented by the wavy line between two magnon bands $\hbar\omega_{1}$ and $\hbar\omega_{2}$ in Fig. 1(b). By developing a quantum kinetic theory for electrical spin injection, dephasing, and transport in noncollinear AFMs, we show that a net spin-flip process of electrons creates definite magnon correlations, illustrated in Fig. 1(b) with canted AFMs as a prototypical example. The injected magnon correlation experiences dephasing due to free-induction decay, governed by the inhomogeneous broadening [48, 46, 47] of oscillation frequencies $|\omega_{1}({\bf k})-\omega_{2}({\bf k})|$ of wave vectors ${\bf k}$, and backaction to spin-flips of electrons. This then provides two routes to control the magnon-correlation transport away from the injector: i) the magnetic field causes the spin spatial oscillation in the transport that is enhanced by the applied magnetic field; ii) a proximity conductor brakes the transport, suppresses spatial oscillation, and thereby acts as a gate. Inducing quantum coherence in incoherent processes may lead to crucial applications and new tunability in future devices.
Spin carried by magnon correlations.—We start from a general Heisenberg Hamiltonian for AFMs $\hat{H}_{m}=(1/2)\sum_{l,l^{\prime}=1}^{\cal N}\sum_{i,j=1}^{N}\sum_{\alpha,\beta}{J}_{ij}^{\alpha\beta}(l,l^{\prime})\hat{S}_{li}^{\alpha}\hat{S}_{l^{\prime}j}^{\beta}$ with noncollinear spins $\hat{\bf S}_{li}$ of the $l$-th ion in the $i$-th magnetic unit cell coupled via the coupling constant $J_{ij}^{\alpha\beta}$. The magnon $\hat{\Psi}_{\bf k}=(\hat{b}_{\xi,{\bf k}},\hat{b}^{\dagger}_{\xi,{\bf-k}})^{T}$ of modes $\xi=\{{1,\cdots,{\cal N}}\}$ and wave vector ${\bf k}$ is found by Bogoliubov transformation with matrix ${\cal T}_{\bf k}$ [49], see Supplemental Material (SM) [50]. The $\alpha$-component of the total spin $\hat{S}_{\alpha}=(1/2)\sum_{\bf k}\hat{\Psi}^{\dagger}_{\bf k}{\cal S}_{\alpha}({\bf k})\hat{\Psi}_{\bf k}$ carried by magnons defines the magnon spin via a matrix
@@
${\cal S}_{\alpha}({\bf k})=-{\cal T}^{\dagger}_{\bf k}\left[I_{{\cal N}\times{\cal N}}\otimes{\rm diag}(\eta^{\alpha}_{1},\cdots,\eta^{\alpha}_{\cal N})\right]{\cal T}_{\bf k},$ (1)
in which $\eta^{\alpha}_{l=\{1,\cdots,{\cal N}\}}$ is the projection of the $l$-th spin ${\bf S}_{l}$ along the lab $\alpha$-axis. The diagonal elements of ${\cal S}_{\alpha}({\bf k})$ are the magnon spin/polarization defined in the classical description [27, 51]. The off-diagonal elements are the magnon correlations; its physical meaning may be fully appreciated in the spin current $\hat{\bf J}_{\alpha}({\bf r})$ and torque $\hat{T}_{\alpha}({\bf r})$ with $\alpha$ denoting the spin polarization. The rate of change of spin density $\hat{S}_{\alpha}({\bf r})$ coverts to the spin current according to $\partial_{t}\hat{S}_{\alpha}({\bf r})=(1/i\hbar)[\hat{S}_{\alpha}({\bf r}),\hat{H}_{m}]=-\nabla\cdot\hat{\bf J}_{\alpha}({\bf r})+\hat{T}_{\alpha}({\bf r})$. When flowing uniformly, the spin current and torque read [52, 53, 50]
@@
$\hat{\bf J}_{0\alpha}$ $=\frac{1}{4V}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\boldsymbol{\cal V}}_{\bf k}+{\boldsymbol{\cal V}}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}},$ (2a)
@@
$\hat{T}_{0\alpha}$ $=\frac{1}{2V}\frac{1}{i\hbar}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\cal E}_{\bf k}-{\cal E}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}},$ (2b)
expressed by the spin matrix (1), mode energies ${\cal E}_{\bf k}={\rm diag}(\hbar\omega_{\xi,\bf k},\hbar\omega_{\xi,\bf-k})$, metric $\sigma_{3}={\rm diag}(I_{{\cal N}\times{\cal N}},-I_{{\cal N}\times{\cal N}})$, and magnon group velocities ${\boldsymbol{\cal V}}_{{\bf k}}=(1/\hbar)\partial{\cal E}_{\bf k}/\partial{\bf k}$.

### v2

@@ Abstract
Thermal and electrical injection and transport of magnon spins in magnetic insulators is conventionally understood by the non-equilibrium population of magnons. However, this view is challenged by several recent experiments in noncollinear antiferromagnets, which urge a thorough theoretical investigation at the fundamental level. We find that the magnon spin in noncollinear magnets is described by a matrix, so even when the diagonal terms—spins carried by population—vanish, the off-diagonal correlations transmit magnon spins. Our quantum theory shows that a net spin-flip of electrons in adjacent conductors creates quantum coherence between magnon states, which transports magnon spins in canted antiferromagnets, even without a definite phase difference between magnon modes in the incoherent process. It reveals that the pumped magnon correlation is not conserved due to an intrinsic spin torque, which causes dephasing and strong spatial spin oscillations during transport; both are enhanced by magnetic fields. Spin transfer to proximity conductors can cause extrinsic dephasing, which suppresses spin oscillations and thereby gates spin transport.
I Introduction
Antiferromagnets (AFMs) hold significant promise for spintronic applications due to their terahertz-frequency dynamics, immunity to external magnetic disorder, and ultrafast responses [1, 2, 3, 4, 5]. Spin angular momentum in collinear easy-axis AFMs is carried by circularly polarized magnon eigenmodes, which shows low dissipation in insulators with efficient long-range spin transport [6, 8, 9, 7, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 21, 22, 20, 23, 24]. In contrast, magnon modes in easy-plane AFMs are linearly polarized and are believed to carry no spins [3, 25, 26, 27] and the modes in canted AFMs have no spin projection along the Néel vector; both appear to be less likely to support spin currents [3, 16, 28, 29]. Experiments on canted easy-plane hematite ($\alpha$-Fe2O3) deny this intuition by observing efficient long-range magnon spin transport and a field-tunable Hanle effect [30, 31, 32]. The observation is interpreted as an interference of two linearly polarized modes, occasionally mimicking a circular polarization or pseudospin [5, 33, 34, 35].
@@ I Introduction
In this work, we demonstrate in the quantum description that the pseudospin in the classical description is a “real spin” carried by magnons by showing that the magnon spin is described by a matrix. When the spin carried by magnon population vanishes due to specific interactions, as shown in Fig. 1(a), the spin is solely carried by correlation $\cal A$}, represented by the wavy line between two magnon bands $\hbar\omega_{1}$ and $\hbar\omega_{2}$ in Fig. 1(b). By developing a quantum kinetic theory for electrical spin injection, dephasing, and transport in noncollinear AFMs, we show that a net spin-flip process of electrons creates definite magnon correlations, illustrated in Fig. 1(b) with canted AFMs as a prototypical example. The injected magnon correlation experiences dephasing due to free-induction decay, governed by the inhomogeneous broadening [48, 46, 47] of oscillation frequencies $|\omega_{1}({\bf k})-\omega_{2}({\bf k})|$ of wave vectors $\bf k$}, and backaction to spin-flips of electrons. This then provides two routes to control the magnon-correlation transport away from the injector: i) the magnetic field causes the spin spatial oscillation in the transport that is enhanced by the applied magnetic field; ii) a proximity conductor brakes the transport, suppresses spatial oscillation, and thereby acts as a gate. Inducing quantum coherence in incoherent processes may lead to crucial applications and new tunability in future magnonic devices.
This article is organized as follows. In Sec. II, we build up the description of magnon spin by a matrix in general magnetic structures (Sec. II.1) and a two-sublattice model of canted AFM (Sec. II.2) as a prototypical example. In Sec. III, we calculate the electrical injection and dephasing of magnon correlation by setting up the quantum kinetic equation. Two functionalities for the spin transport by magnon correlation, including the Hanle effect and gating effect by metals, are addressed in Sec. IV. We summarize our results and give an outlook in Sec. V.
II Spin carried by magnon correlations
II.1 General magnetic structures
We start from a general Heisenberg Hamiltonian for noncollinear magnets
$\hat{H}_{m}=(1/2)\sum_{l,l^{\prime}=1}^{\cal N}\sum_{i,j=1}^{N}\sum_{\alpha,\beta}{J}_{ij}^{\alpha\beta}(l,l^{\prime})\hat{S}_{li}^{\alpha}\hat{S}_{l^{\prime}j}^{\beta}$ (1)
with noncollinear spins $\hat{\bf S}_{li}$ of the $l$-th ion in the $i$-th magnetic unit cell coupled via the coupling constant $J_{ij}^{\alpha\beta}$. The equilibrium configuration of $\cal N$} spins with their directions ${\boldsymbol{\eta}}_{1,\cdots,{\cal N}}$ is illustrated in Fig. 2. For the $l$-th magnetic ions in one magnetic unit cell, we define a local $\{\hat{\bf x}^{\prime}_{l},\hat{\bf y}^{\prime}_{l},\hat{\bf z}^{\prime}_{l}\}$-frame with $\hat{\bf z}^{\prime}_{l}={\boldsymbol{\eta}}_{l}$ along its ground-state spin, so the $\cal N$} sublattices in one magnetic unit cell with spins along different directions imply $\cal N$} local frames.
Figure 2: A general magnetic structure with $\cal N$} noncollinear spins in one unit cell. ${\boldsymbol{\eta}}_{l=\{1,2\cdots,{\cal N}\}}$ represents their directions in the ground state.
In the local frames, we represent the spin operator $\hat{\bf S}_{li}$ by the bosonic operators $\hat{a}_{li}$ via the Holstein-Primakoff transformation. In the linear-response regime, with $|\hat{\bf S}_{li}|=S$,
$\hat{S}_{li}^{x^{\prime}}$ $=\frac{\sqrt{2S}}{2}\left(\hat{a}_{li}+\hat{a}_{li}^{\dagger}\right),$ (2)
$\hat{S}_{li}^{y^{\prime}}$ $=\frac{\sqrt{2S}}{2i}\left(\hat{a}_{li}-\hat{a}_{li}^{\dagger}\right),$
$\hat{S}_{li}^{z^{\prime}}$ $=S-\hat{a}_{li}^{\dagger}\hat{a}_{li}.$
The local $\{\hat{\bf x}^{\prime},\hat{\bf y}^{\prime},\hat{\bf z}^{\prime}\}$-frame is transformed to the lab $\{\hat{\bf x},\hat{\bf y},\hat{\bf z}\}-$frame via the rotation matrices $\left(\hat{\bf x}^{\prime}_{l},\hat{\bf y}^{\prime}_{l},\hat{\bf z}^{\prime}_{l}\right)$ according to
$\left(\begin{array}[]{c}\hat{S}_{li}^{x}\\ \hat{S}_{li}^{y}\\ \hat{S}_{li}^{z}\end{array}\right)=\left({\bf x}^{\prime}_{l},{\bf y}^{\prime}_{l},{\bf z}^{\prime}_{l}\right)\left(\begin{array}[]{c}\hat{S}_{li}^{x^{\prime}}\\ \hat{S}_{li}^{y^{\prime}}\\ \hat{S}_{li}^{z^{\prime}}\end{array}\right),$ (3)
such that the spin operators in the lab coordinate frame read
$\hat{\bf S}_{li}=\frac{\sqrt{2S}}{2}{\boldsymbol{\xi}}_{l}^{\ast}\hat{a}_{li}+\frac{\sqrt{2S}}{2}{\boldsymbol{\xi}}_{l}\hat{a}^{\dagger}_{li}+{\boldsymbol{\eta}}_{l}(S-\hat{a}^{\dagger}_{li}\hat{a}_{li}).$ (4)
Here, ${\boldsymbol{\xi}}_{l}=\hat{\bf x}^{\prime}_{l}+i\hat{\bf y}^{\prime}_{l}$ represents the ellipticity of the spin precession and ${\boldsymbol{\eta}}_{l}=\hat{\bf z}^{\prime}_{l}$ is the direction of ground-state spins.
After the Fourier transformation over the magnetic unit cells $\{i,j\}$, we obtain the Hamiltonian
$\hat{H}_{m}=\frac{1}{2}\sum_{\bf k}\hat{X}_{\bf k}^{\dagger}{\cal H}_{\bf k}\hat{X}_{\bf k},$ (5)
in which $\hat{X}_{\bf k}=\left(\hat{a}_{l,{\bf k}},\\ \hat{a}^{\dagger}_{l,{\bf-k}}\right)^{T}$ with $\{l=1,...,{\cal N}\}$ and ${\cal H}_{\bf k}$ is the Hamiltonian matrix. In terms of a paraunitary matrix ${\cal T}_{\bf k}$ with ${\cal T}_{\bf k}\sigma_{3}{\cal T}_{\bf k}^{\dagger}=\sigma_{3}$ or ${\cal T}_{\bf k}^{\dagger}\sigma_{3}{\cal T}_{\bf k}=\sigma_{3}$, where $\sigma_{3}=\begin{pmatrix}I_{{\cal N}\times{\cal N}}&0\\ 0&-I_{{\cal N}\times{\cal N}}\end{pmatrix}$ is the metric matrix, the mode energies ${\cal E}_{\bf k}\equiv{\rm diag}(\hbar\omega_{1,\bf k},...,\hbar\omega_{{\cal N},\bf k},\hbar\omega_{1,\bf-k},...,\hbar\omega_{{\cal N},\bf-k})={\cal T}_{\bf k}^{\dagger}{\cal H}_{\bf k}{\cal T}_{\bf k}$ [49]. The eigenvalue problem is solved for the matrix $\sigma_{3}{\cal H}_{\bf k}$ and the eigenvectors $u_{i}$ of $\sigma_{3}{\cal H}_{\bf k}$ can be arranged to express the transformation matrix ${\cal T}_{\bf k}$. Accordingly,
$\hat{H}_{m}=\frac{1}{2}\sum_{\bf k}\hat{\Psi}_{\bf k}^{\dagger}({\cal T}^{\dagger}_{\bf k}{\cal H}_{\bf k}{\cal T}_{\bf k})\hat{\Psi}_{\bf k},$ (6)
in which $\hat{\Psi}_{\bf k}=\left(\hat{b}_{\xi,{\bf k}},\\ \hat{b}^{\dagger}_{\xi,{\bf-k}}\right)^{T}$ defines the magnon operators of modes $\xi=\{1,...,{\cal N}\}$ via $\hat{X}_{\bf k}={\cal T}_{\bf k}\hat{\Psi}_{\bf k}$.
The spin matrix of the magnon is found by rewriting Eq. (4) in the basis of the magnon operators using the Bogoliubov transformation matrix ${\cal T}_{\bf k}$. The $\alpha$-component of the total-spin operator
${\color[rgb]{0,0,1}\hat{S}^{\rm tot}_{\alpha}}$ $=-\frac{1}{2}\sum_{\bf k}\hat{X}^{\dagger}_{\bf k}\Big[I_{2\times 2}\otimes{\rm diag}(\eta^{\alpha}_{1},...,\eta^{\alpha}_{\cal N})\Big]\hat{X}_{\bf k}$
$+{\rm zeroth~and~first~order~terms~of~}\{\hat{a},\hat{a}^{\dagger}\},$ (7)
in which $\eta^{\alpha}_{l=\{1,\cdots,{\cal N}\}}$ is the projection of ${\boldsymbol{\eta}}_{l}$ along the $\alpha$-axis of the lab frame. The zeroth and first orders of the bosonic operators $\{\hat{a},\hat{a}^{\dagger}\}$ have no contribution to the magnon spin. Through the paraunitary matrix ${\cal T}_{\bf k}$,
$\hat{S}_{\alpha}=(1/2)\sum_{\bf k}\hat{\Psi}^{\dagger}_{\bf k}{\cal S}_{\alpha}({\bf k})\hat{\Psi}_{\bf k}$ (8)
is transformed into the magnon basis with the matrix
${\cal S}_{\alpha}({\bf k})$ $=-{\cal T}^{\dagger}_{\bf k}\left[I_{{2}\times{2}}\otimes{\rm diag}(\eta^{\alpha}_{1},...,\eta^{\alpha}_{\cal N})\right]{\cal T}_{\bf k}.$ (9)
The diagonal elements of ${\cal S}_{\alpha}({\bf k})$ are the magnon spin/polarization defined in the classical description [27, 50]. As illustrated in Fig. 3, the $2{\cal N}\times 2{\cal N}$ spin matrix contains four ${\cal N}\times{\cal N}$ blocks, represented, respectively, by $(\hat{b}^{\dagger}\hat{b})_{{\cal N}\times{\cal N}}$, $(\hat{b}^{\dagger}\hat{b}^{\dagger})_{{\cal N}\times{\cal N}}$, $(\hat{b}\hat{b})_{{\cal N}\times{\cal N}}$, and $(\hat{b}\hat{b}^{\dagger})_{{\cal N}\times{\cal N}}$ blocks. The $(\hat{b}^{\dagger}\hat{b}^{\dagger})_{{\cal N}\times{\cal N}}$ and $(\hat{b}\hat{b})_{{\cal N}\times{\cal N}}$ blocks are magnon spins carried by two creation or annihilation operators. In the $(\hat{b}^{\dagger}\hat{b})_{{\cal N}\times{\cal N}}$ block, the off-diagonal elements are the magnon correlations; its physical meaning may be fully appreciated in the spin current $\hat{\bf J}_{\alpha}({\bf r})$ and torque $\hat{T}_{\alpha}({\bf r})$ with $\alpha$ denoting the spin polarization.
Figure 3: The structure of the spin matrix.
Equation (9) is defined for any magnetic configuration. There have been specific cases in the literature where the description of the magnon spin matrix is provided. For example, Refs. [51] and [27] defines the spin $S^{\alpha}_{m,\bf k}=\langle m,{\bf k}|\hat{S}_{\alpha}|m,{\bf k}\rangle-\langle 0|\hat{S}_{\alpha}|0\rangle$ carried by the magnon mode $m=\{1,...,{\cal N}\}$, which only addresses the diagonal terms ${\cal S}_{\alpha}({\bf k})]_{mm}$ of the spin matrix, as shown in Fig. 3. These spins are carried by the magnon population that is responsible for the spin Seebeck effect. On the other hand, Ref. [52] investigated the excitation of Brillouin zone-edge magnon-pairs ($\hat{b}^{\dagger}_{l}\hat{b}^{\dagger}_{l^{\prime}}$ with $l\neq l^{\prime}$) using optical pulses. In their subsequent work, Ref. [53] discussed ultrafast, optically-controlled spin currents carried by such magnon pairs in collinear antiferromagnets.
The matrix description represents the real spin in the sense that it originates from the magnetic moments of every ion, as in Fig. 2. For a magnetic unit cell with $\cal N$} sublattices, the creation of a Holstein-Primakoff boson decreases the spin $\hbar$ along the magnetic moment of that ion. The magnon operator connects to the Holstein-Primakoff boson via the Bogoliubov transformation, which then carries the real spin of the system. This microscopic perspective distinguishes itself from other discussions of angular momentum. For instance, Ref. [54] focuses on ferromagnets, which include the orbital angular momentum that is absent in our canted antiferromagnets. Reference [55] addresses the pseudo-orbital angular momentum of phonons and magnons, while we focus on spin angular momentum. Reference [33] constructs the pseudospin $(\tilde{L}_{x},\tilde{L}_{y},\tilde{L}_{z})$, which is only a convenience to expand a $2\times 2$ Hamiltonian matrix of an antiferromagnet by the Holstein-Primakoff bosons. In such a peusospin description, only $\tilde{L}_{z}$ represents the real spin of the magnon along the Néel vector, while $\tilde{L}_{x}$ and $\tilde{L}_{y}$ do not correspond to measurable spins.
The off-diagonal magnon correlations contribute to the spin matrix when the associated spin component is not conserved, i.e., when $\hat{S}_{\alpha},\hat{H}_{m}]\neq 0$ due to broken spin rotational symmetry. By substituting Eqs. (6) and (8), we find
$[\hat{S}_{\alpha},\hat{H}_{m}]=\frac{1}{2}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{\bf k}[{\cal S}_{\alpha}({\bf k})\sigma_{3}{\cal E}_{\bf k}-{\cal E}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})]\hat{\Psi}_{\bf k}.$ (10)
Accordingly, when $\hat{S}_{\alpha},\hat{H}_{m}]=0$ and ${\cal E}_{\bf k}$ is non-degenerate, the spin matrix ${\cal S}_{\alpha}({\bf k})$ should be diagonal, i.e., the magnon correlation contributes no spin operator. In other words, when $\hat{S}_{\alpha},\hat{H}_{m}]\neq 0$, the off-diagonal components in the spin matrix exist and contribute to the magnon spin. In collinear systems, such as conventional ferromagnets and easy-axis antiferromagnets, where $U(1)$ spin symmetry around the easy-axis is preserved, the spin matrix is diagonal, such that the diagonal terms solely describe spin. Conversely, in noncollinear and canted antiferromagnets, the $U(1)$ symmetry can be broken, rendering the off-diagonal correlation terms a contribution to the spin.
The spin current and spin torque densities of magnons are derived from the continuity equation of the spin-density operator
$\hat{S}_{\alpha}({\bf r})$ $=-(1/2)\hat{X}^{\dagger}({\bf r})\left[I_{{2}\times{2}}\otimes{\rm diag}(\eta^{\alpha}_{1},...,\eta^{\alpha}_{\cal N})\right]\hat{X}({\bf r})$
$=\frac{1}{V}\sum_{\bf q}\hat{D}_{\alpha}({\bf q})e^{i{\bf q}\cdot{\bf r}}.$ (11)
The Fourier component of the spin-density operator reads
@@ II.1 General magnetic structures
$\hat{D}_{\alpha}({\bf q})$ $=-\frac{1}{2}\sum_{\bf k}\hat{X}^{\dagger}_{\bf k}\left[I_{{2}\times{2}}\otimes{\rm diag}(\eta^{\alpha}_{1},...,\eta^{\alpha}_{\cal N})\right]\hat{X}_{\bf k+q}$
$=\frac{1}{2}\sum_{\bf k}\hat{\Psi}^{\dagger}_{\bf k}{\cal D}_{\alpha}({\bf k},{\bf q})\hat{\Psi}_{\bf k+q},$ (12)
where the matrix
${\cal D}_{\alpha}({\bf k},{\bf q})\equiv-{\cal T}^{\dagger}_{\bf k}\left[I_{{2}\times{2}}\otimes{\rm diag}(\eta^{\alpha}_{1},...,\eta^{\alpha}_{\cal N})\right]{\cal T}_{\bf k+q}.$
In the continuity equation
$\frac{\partial\hat{S}_{\alpha}({\bf r})}{\partial t}=-\nabla\cdot\hat{\bf J}_{\alpha}({\bf r})+\hat{T}_{\alpha}({\bf r}),$ (13)
$\hat{\bf J}_{\alpha}({\bf r})$ and $\hat{T}_{\alpha}({\bf r})$ represent the spin-current and spin-torque densities. According to $\nabla\cdot\hat{\bf J}_{\alpha}({\bf r})=(1/V)\sum_{\bf q}i{\bf q}\cdot\hat{\bf J}_{\alpha}({\bf q})e^{i{\bf q}\cdot{\bf r}}$, Eq. (13) implies
$\frac{\partial\hat{D}_{\alpha}({\bf q})}{\partial t}=-i{\bf q}\cdot\hat{\bf J}_{\alpha}({\bf q})+\hat{T}_{\alpha}({\bf q}).$ (14)
On the other hand, by the Heisenberg equation of motion,
$\frac{\partial\hat{D}_{\alpha}({\bf q})}{\partial t}=\frac{1}{i\hbar}\left(\left[\hat{D}_{\alpha}({\bf q}),\hat{H}_{m}\right]+\left[\hat{D}_{\alpha}({\bf q}),\hat{H}_{\rm int}\right]\right),$
where $\hat{H}_{\rm int}$ originates from the interplay with environment. The second term contributes an external spin torque $\hat{T}_{\rm ext}$ induced by interactions with the environment. We are concerned with the intrinsic contribution, which leads to the spin-current operator and the intrinsic spin torque $\hat{T}_{\rm in}$.
Based on the commutators between the magnon operators $\hat{b}_{\xi,{\bf k}},\hat{b}^{\dagger}_{\xi^{\prime},\bf k^{\prime}}]=\delta_{\xi\xi^{\prime}}\delta_{\bf kk^{\prime}}$ and $\hat{b}_{\xi,{\bf k}},\hat{b}_{\xi^{\prime},\bf k^{\prime}}]=0$, the field operators $\hat{\Psi}_{\bf k}=\left(\hat{b}_{1,{\bf k}},\\ \hat{b}_{2,{\bf k}},\\ \hat{b}^{\dagger}_{1,{\bf-k}},\\ \hat{b}^{\dagger}_{2,{\bf-k}}\right)^{T}$ obey $\hat{\Psi}_{n,{\bf k}},\hat{\Psi}^{\dagger}_{m,{\bf k^{\prime}}}]=(\sigma_{3})_{nm}\delta_{\bf kk^{\prime}}$, $\hat{\Psi}_{n,{\bf k}},\hat{\Psi}_{m,{\bf k^{\prime}}}]=(i\sigma_{2})_{nm}\delta_{{\bf-k},{\bf k^{\prime}}}$, and $\hat{\Psi}^{\dagger}_{n,{\bf k}},\hat{\Psi}^{\dagger}_{m,{\bf k^{\prime}}}]=(-i\sigma_{2})_{nm}\delta_{{\bf-k},{\bf k^{\prime}}}$, where $\sigma_{2}=\begin{pmatrix}0&-iI_{{\cal N}\times{\cal N}}\\ iI_{{\cal N}\times{\cal N}}&0\end{pmatrix}$. By combining Eqs. (6) and (12), and with the diagonal matrix ${\cal E}_{\bf k}={\cal T}_{\bf k}^{\dagger}{\cal H}_{\bf k}{\cal T}_{\bf k}$, we find
$[\hat{D}_{\alpha}({\bf q}),\hat{H}_{m}]$
$=\frac{1}{2}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{\bf k}\left({\cal D}_{\alpha}({\bf k},{\bf q})\sigma_{3}{\cal E}_{\bf k+q}-{\cal E}_{\bf k}\sigma_{3}{\cal D}_{\alpha}({\bf k},{\bf q})\right)\hat{\Psi}_{\bf k+q}.$
When the external driven field varies slowly in the real space, we participate only the excitation of small wave vector $\bf q$} (long wavelength) is involved. By shifting ${\bf k}\rightarrow{\bf k}-{\bf q}/2$, to the leading order of $\bf q$}, we obtain the spin-current density operator [56],
$\hat{\bf J}_{\alpha}({\bf q})$ $=\frac{1}{4}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}-{\bf q}/2}({\cal D}_{\alpha}({\bf k}-{\bf q}/2,{\bf q})\sigma_{3}{\boldsymbol{\cal V}}_{\bf k}$
$+{\boldsymbol{\cal V}}_{\bf k}\sigma_{3}{\cal D}_{\alpha}({\bf k}-{\bf q}/2,{\bf q}))\hat{\Psi}_{{\bf k}+{\bf q}/2},$ (16)
and the intrinsic spin-torque density operator
$\hat{T}^{\alpha}_{{\rm in}}({\bf q})$ $=\frac{1}{2i\hbar}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}-{\bf q}/2}({\cal D}_{\alpha}({\bf k}-{\bf q}/2,{\bf q})\sigma_{3}{\cal E}_{\bf k}$
$-{\cal E}_{\bf k}\sigma_{3}{\cal D}_{\alpha}({\bf k}-{\bf q}/2,{\bf q}))\hat{\Psi}_{{\bf k}+{\bf q}/2},$ (17)
in which magnon group velocities ${\boldsymbol{\cal V}}_{{\bf k}}=(1/\hbar)\partial{\cal E}_{\bf k}/\partial{\bf k}$. The non-vanishing spin torque density indicates the non-conservation of the injected spin.
When flowing uniformly, ${\bf q}=0$ in Eq. (16), such that the spin current and torque read [57, 56]
@@
$\hskip-8.5359pt\hat{\bf J}_{0\alpha}$ $=\frac{1}{4V}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\boldsymbol{\cal V}}_{\bf k}+{\boldsymbol{\cal V}}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}},$ (18a)
@@
$\hskip-8.5359pt\hat{T}_{0\alpha}$ $=\frac{1}{2V}\frac{1}{i\hbar}\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}\left({\cal S}_{\alpha}({\bf k})\sigma_{3}{\cal E}_{\bf k}-{\cal E}_{\bf k}\sigma_{3}{\cal S}_{\alpha}({\bf k})\right)\hat{\Psi}_{{\bf k}}.$ (18b)
II.2 Two-sublattice model of AFM
The magnon polarization/spin in the classical description is often quenched by specific interactions since they hybridize two circularly polarized modes, such as local Dzyaloshinskii-Moriya interaction [27, 58, 59] or transverse magnetic field $\bf H$}, rendering AFMs canted. Using a two-sublattice AFM as a prototypical model, we illustrate the spin matrix and its physical meaning.
As depicted in Fig. 4(a), two nearest-neighboring spins couple via the antiferromagnetic exchange interaction $J>0$ and are aligned along the Néel vector $\hat{\bf z}-$axis by the easy-axis anisotropy $K_{z}<0$, governed by the Hamiltonian
$\hat{H}_{m}$ $=J\sum_{\langle i,j\rangle}\hat{\bf S}_{1i}\cdot\hat{\bf S}_{2j}+K_{z}\sum_{l}^{2}\sum_{i}(\hat{S}_{li}^{z})^{2}$
$-|\gamma|\hbar\mu_{0}H\sum_{l}^{2}\sum_{i}\hat{S}^{x}_{li},$ (19)
where $\langle\cdots\rangle$ represents the summation over the nearest-neighboring sites $\{i,j\}$ and $\gamma<0$ is the gyromagnetic ratio of electrons. A magnetic field $\bf H$} along the $\hat{\bf x}-$direction stabilizes a canted spin configuration with $\hat{\bf S}_{1}=S(\sin\theta,0,\cos\theta)$ and $\hat{\bf S}_{2}=S(\sin\theta,0,-\cos\theta)$ not parallel in one magnetic unit cell by a canted angle $\theta={\rm arcsin}\left(|\gamma|\hbar\mu_{0}H/(8JS-2K_{z}S)\right)$. In the lab coordinate, ${\bf S}_{1,2}$ are along ${\boldsymbol{\eta}}_{1(2)}=(\sin\theta,0,\pm\cos\theta)^{T}$. There are two modes with the field operator of magnons $\hat{\Psi}_{{\bf k}}=(\hat{b}_{1,{\bf k}},\hat{b}_{2,{\bf k}},\hat{b}^{\dagger}_{1,-{\bf k}},\hat{b}^{\dagger}_{2,-{\bf k}})^{T}$.
According to Eq. (4), the Hamiltonian $\hat{H}_{m}=(1/2)\sum_{\bf k}\hat{X}^{\dagger}_{\bf k}{\cal H}_{\bf k}\hat{X}_{\bf k}$ in terms of the bosonic operators $\hat{X}_{\bf k}=(\hat{a}_{1,{\bf k}},\hat{a}_{2,{\bf k}},\hat{a}^{\dagger}_{1,-{\bf k}},\hat{a}^{\dagger}_{2,-{\bf k}})^{T}$, where the Hamiltonian matrix
${\cal H}({\bf k})=S\begin{pmatrix}A&C_{+}({\bf k})&B&C_{-}({\bf k})\\ C_{+}({\bf k})&A&C_{-}({\bf k})&B\\ B&C_{-}({\bf k})&A&C_{+}({\bf k})\\ C_{-}({\bf k})&B&C_{+}({\bf k})&A\end{pmatrix}$. (20)